# Exact Simulation of Longitudinal Data from Marginal Structural Models

### v2

@@ 4.3 Simulation Algorithm
Extending the simulation algorithm to include a latent process $\{H_{k}\}_{k\in\mathcal{K}}$, or multiple time-varying confounders is straightforward. The key modification lies in the outcome simulation: the marginal quantile of $Y$ is obtained by sequentially removing the conditioning on all the time-dependent confounders at each step, through inverse $h$-functions. We provide an illustrative example including a latent variable $H_{k}$ in Appendix A.
@@ 5 Simulating Survival Outcomes
  • Continuous-time hazard MSM, where we introduce an auxiliary continuous variable, $\tilde{Y}_{k+1}=T-k\,|\,T>k$, representing the incremental failure time given survival at the start of the $k$th interval. We model incremental survival time $\tilde{Y}_{k+1}$ and truncated it to obtain the observed binary survival outcome $Y_{k+1}$. Given survival at $k$, the time to failure is $T\,|\,T>k=k+\tilde{Y}_{k+1}$. Examples of marginal models for $\tilde{Y}_{k+1}$ include:
    • Cox proportional hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)\exp\{g(\bar{a}_{k},z;\beta)\}$,
    • Additive hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)+g(\bar{a}_{k},z;\gamma)$.
@@ 5.1 Simulation algorithm
Unlike the derivation of outcome simulation for the non-survival case in Section 4.2, all quantities involved in the copulas must now be conditioned on survival up to this point. Specifically, when simulating the incremental failure time, $\tilde{y}_{k+1}$, given $Y_{k}=0$, (7) becomes:
$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
$=\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0),$
@@
It is important to note the distinction between the subscripts $j$ and $k$ in the product. To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, which corresponds to the $k$th panel in Figure 2. Each copula in (10) associates $\tilde{y}_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,\ldots,k$.
In the non-survival case discussed in Section 4.2, the quantiles $F(l_{j}\,|\,\bar{l}_{j-1},\bar{a}_{j-1})$, for $j=0,\ldots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation. However, this convenience does not extend when the quantiles are conditioned on survival at later time points. For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim\operatorname{N}(0,1)$. Conditioning on survival after 12 months, however, alters the distribution. The baseline health of the survivors will almost certainly not be representative of the initial population, meaning that the distribution is no longer standard normal. This means that the conditional quantiles of $L_{j}$ that we need are not directly available and require additional computation. Fortunately, employing copulas to define the dependence structure offers an efficient method to derive these conditional distributions, eliminating the need for resource-intensive techniques such as Monte Carlo approximation or numerical integration.
The proposed simulation algorithm follows an iterative structure and is formally presented and validated using an inductive approach.
@@
$F_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}$ $F_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}$ $F_{\tilde{Y}_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}.$
Additionally, we define $k+1$ bivariate copulas $C_{\tilde{Y}_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$. These copulas captures the dependence between $\tilde{Y}_{k+1}$ and each preceding time-varying covariate. Draw $w_{k,1}$, $w_{k,2}$, $w_{k,3}\sim\operatorname{U}(0,1)$ and compute:
@@
to obtain a sample $(l_{k},a_{k})$. We initialize by setting $F(\tilde{y}_{k+1}\,|\,\bar{l}_{k},do(\bar{a}_{k}),y_{k}=0)=w_{k,3}$. Then, for $j=k-1,\ldots,0$, we iteratively apply the transformations:
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
$\qquad\quad F(\tilde{y}_{k}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0);\Theta_{{k-1},j}\big{\}}$ (12)
$F(\tilde{y}_{k+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j+1},do(\bar{a}_{k}),y_{k}=0),$
@@
By induction, the quantiles in the $h$-function in (12) are available from the previous time step $k-1$. We prove this claim as Proposition 5.1.
When the recursion reaches $j=0$, we obtain the quantile $F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0)$. Applying the $h$-function one final time, we set:
$u_{k+1}=F(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}))$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0),$
@@
A sample of $\tilde{y}_{k+1}$ is then obtained via inversion:
$\tilde{y}_{k+1}=F_{Y_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}^{-1}(u_{k+1}),$ (14)
with the corresponding binary outcome $y_{k+1}=I(\tilde{y}_{k+1}<1)$. This completes the $k$th time step, yielding the sample $(l_{k},a_{k},\tilde{y}_{k+1},y_{k+1})$. If $y_{k+1}=1$, the simulation stops; otherwise, proceed to $k+1$.
Crucially, our method avoids computing the full conditional distribution of $\tilde{Y}_{k+1}\mid\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0$, but rather requiring just a quantile corresponding to our simulated value; this is central to the efficiency of our approach. Equation (12) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$. In (13), we recursively apply $k+1$ inverse $h$-functions to obtain the marginal quantile of $F(\tilde{Y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0)$. Finally, we simulate $\tilde{Y}_{k+1}$ from the target marginal structural model via inversion. In total, this time step requires $42k+$ 4 functions, including the simulation of $L_{k}$ and $A_{k}$.
@@
The proof is given in Appendix B.
@@
To simulate a binary survival outcome directly from a discrete marginal structural model, such as that defined in (9), we adopt an approach analogous to Fan et al. (2017). Specifically, we assume that there exists a continuous latent random variable, $S_{k+1}$, whose distribution under the intervention $do(\bar{A}_{k})$ is $F_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}$, such that the binary survival indicator is obtained by dichotomizing $S_{k+1}$ at a constant threshold $C$. This formulation allows dependencies between $Y_{k+1}$ and past covariates to be modelled via the latent variable $S_{k+1}$, and the simulation procedure follows the algorithm described above, replacing $\tilde{Y}_{k+1}$ by $S_{k+1}$. Given the marginal quantile $u_{k+1}$, inversion yields $s_{k+1}=F^{-1}_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)$ Finally, the binary outcome is obtained as $y_{k+1}=I(s_{k+1}>C)$.
5.2 Extensions and computational complexity
@@
In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(2k+4)=K^{2}+3K$ calculations. This implies that the computational complexity grows quadratically with $K$. However, if an individual experiences failure at any time step, the simulation terminates early, meaning that this complexity represents the maximum possible computational cost per individual.
A key extension involves simulating multiple outcomes, and the approach depends on whether the additional outcomes pertain to survival or non-survival events. For non-survival outcomes simulated from marginal structural models—as opposed to conditional distributions—the number of calculations to derive $F(\bm{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y^{\triangle}_{k}=0)$ increases proportionally with the number of outcomes, $m$. Here, $\bm{Y}_{k+1}$ is an $m$-dimensional vector, and $Y_{k}^{\triangle}$ denotes the primary survival outcome, with all other outcomes assumed to be non-survival. To incorporate dependence between outcomes given past variables, we must specify the joint distribution $\bm{Y}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k}$ and simulate from this using a pair-copula construction. The total number of calculations is $\sum_{k=0}^{K-1}\big{\{}(k+3)m+k+1\big{\}}=(K+5)Km/2+K(K+1)/2$, which scales linearly with $m$.
For competing risks, such as death from other causes, additional calculations are needed to determine the distribution of $L$ given survival from all risks, as well as its distribution conditioned on the occurrence of each specific risk. Consequently, the computational cost scales proportionally with $m$. Furthermore, iterative evaluations of inverse $h$-functions are required to obtain marginal interventional quantiles for each risk. The total number of calculations is given by $\sum_{k=0}^{K-1}(2mk+3m+1)=(K^{2}+2K)m+K$ which grows linearly with $m$. The simulation algorithm for a setting with a single competing risk is described in Appendix C.

### v3

@@ 4.3 Simulation Algorithm
Extending the simulation algorithm to incorporate multiple time-varying confounders—whether assumed to be observed or unobserved — is straightforward. The key modification lies in the outcome simulation: the marginal quantile of $Y$ is obtained by sequentially removing the conditioning on all the time-dependent confounders at each step, through inverse $h$-functions. We provide an illustrative example in Appendix A, outlining the simulation procedure for incorporating an additional latent process $\{H_{k}\}_{k\in\mathcal{K}}$.
@@ 5 Simulating Survival Outcomes
  • Continuous-time hazard MSM, where we introduce an auxiliary continuous variable, $\widetilde{Y}_{k+1}=T-k\,|\,T>k$, representing the incremental failure time given survival at the start of the $k$th interval. We model incremental survival time $\widetilde{Y}_{k+1}$ and truncated it to obtain the observed binary survival outcome $Y_{k+1}$. Given survival at $k$, the time to failure is $T\,|\,T>k=k+\widetilde{Y}_{k+1}$. Examples of marginal models for $\widetilde{Y}_{k+1}$ include:
    • Cox proportional hazard:
      $\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)\exp\{g(\bar{a}_{k},z;\beta)\},$
    • Additive hazard:
      $\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)+g(\bar{a}_{k},z;\gamma).$
@@ 5.1 Simulation algorithm
Unlike the derivation of outcome simulation for the non-survival case in Section 4.2, all quantities involved in the copulas must now be conditioned on survival up to this point. Specifically, when simulating the incremental failure time, $\widetilde{y}_{k+1}$, given $Y_{k}=0$, (7) becomes:
$p(\widetilde{y}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},y_{k}=0)=$ $p^{*}(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)\cdot$
$\prod_{j=0}^{k}c^{*}_{\widetilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0}\big{\{}F(\widetilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)),$
@@
It is important to note the distinction between the subscripts $j$ and $k$ in the product. To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, which corresponds to the $k$th panel in Figure 2. Each copula in (11) associates $\widetilde{y}_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,\ldots,k$.
In the non-survival case discussed in Section 4.2, the quantiles $F(l_{j}\,|\,\bar{l}_{j-1},\bar{a}_{j-1})$, for $j=0,\ldots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation. However, this convenience does not extend when the quantiles are conditioned on survival at later time points. For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim\operatorname{N}(0,1)$. Conditioning on survival after 12 months, however, alters the distribution. The baseline health of the survivors will almost certainly not be representative of the initial population, meaning that the distribution is no longer standard normal. This means that the conditional quantiles of $L_{j}$ that we need are not directly available and require additional computation. Fortunately, employing copulas to define the dependence structure offers a solution to deriving these conditional distributions.
Lemma 5.1.
In the setting described above, suppose $Y_{k}=0\Leftrightarrow\widetilde{Y}_{k}>1$, then the conditional distribution $F_{L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$ is given by
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=\frac{v_{k-1,j}-C_{\widetilde{Y}_{k}L_{j}\mid\bar{L}_{j},do(\bar{A}_{k}),Y_{k-1}=0}(v_{k-1,j},q_{k-1,j};\Theta_{k-1,j})}{1-q_{k-1,j}},$ (12)
where $v_{k-1,j}=F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$ and $q_{k-1,j}=F(\widetilde{y}_{k}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$.
The proof is given in Appendix B. Here we emphasize that $F_{L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$ is the “renormalized” distribution in the population who have survived up to time-point $k$. This is distinct from simply conditioning on an individual’s previous incremental failure time, which would give $F_{L_{j}\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),\widetilde{Y}_{k}}$. The former corresponds to the expectation of the latter with respect to $\widetilde{Y}_{k}$ given $\widetilde{Y}_{k}>1$.
The proposed simulation algorithm combines an iterative and recursive structure and is formally presented and validated using an inductive approach.
@@
$F_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}$ $F_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}$ $F_{\widetilde{Y}_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}.$
Additionally, we define $k+1$ bivariate copulas $C_{\widetilde{Y}_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$. These copulas captures the dependence between $\widetilde{Y}_{k+1}$ and each preceding time-varying covariate. Draw $w_{k,1}$, $w_{k,2}$, $w_{k,3}\sim\operatorname{U}(0,1)$ and compute:
@@
to obtain a sample $(l_{k},a_{k})$. We initialize by setting $F(\widetilde{y}_{k+1}\,|\,\bar{l}_{k},do(\bar{a}_{k}),y_{k}=0)=w_{k,3}$. Then, for $j=k-1,\ldots,0$, we iteratively apply the transformations:
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=\frac{v_{k-1,j}-C_{\widetilde{Y}_{k}L_{j}\mid\bar{L}_{j},do(\bar{A}_{k}),Y_{k-1}=0}(v_{k-1,j},q_{k-1,j};\Theta_{k-1,j})}{1-q_{k-1,j}},$ (14)
$F(\widetilde{y}_{k+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ $=h^{-1}\big{\{}F(\widetilde{y}_{k+1}\,|\,\bar{l}_{j+1},do(\bar{a}_{k}),y_{k}=0),$
@@
Expression (14) follows Lemma 5.1, by induction, $v_{k-1,j}$ and $q_{k-1,j}$ are available from the previous time step $k-1$.
When the recursion reaches $j=0$, we obtain the quantile $F(\widetilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0)$. Applying the $h$-function one final time, we set:
$u^{*}_{k+1}=F(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}))$ $=h^{-1}\big{\{}F(\widetilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0),$
@@
A sample of $\widetilde{y}_{k+1}$ is then obtained via inversion:
$\widetilde{y}_{k+1}=F_{Y_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}^{-1}(u^{*}_{k+1}),$
with the corresponding binary outcome $y_{k+1}=I(\widetilde{y}_{k+1}<1)$. If $y_{k+1}=1$, the simulation stops; otherwise, before proceeding to $k+1$, we store and compute the following quantities for the next step:
  • Store $v_{k,j}$ = $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$, for $j=0,\ldots,k$ , which have been computed in the current step via (14).
  • Compute $q_{k,j}=F(\widetilde{y}_{k+1}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$: initialize $q_{k,0}=F(\widetilde{y}_{k+1}=1\,|\,do(\bar{a}_{k}),y_{k}=0)$, then for $j=1,\ldots k$, recursively update:
    $q_{k,j}$ $=F(\widetilde{y}_{k+1}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$
    $=h\big{\{}F(\widetilde{y}_{k+1}=1\,|\,\bar{l}_{j-2},do(\bar{a}_{k}),y_{k}=0),F(l_{j-1}\,|\,\bar{l}_{j-2},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j-1}\big{\}}$
    $=h\big{\{}q_{k,j-1},v_{k,j-1};\Theta_{k,j-1}\big{\}}.$ (16)
This completes the $k$th time step, yielding the sample $(l_{k},a_{k},\widetilde{y}_{k+1},y_{k+1})$, along with the quantities required for the next step. We formalize this claim as Proposition 5.2.
Algorithm 2 Simulation algorithm for an individual $i$ up to the $K$th follow-up
$k\leftarrow 0$
while $k\leq K$ do
     Sample $w_{k,1},w_{k,2},w_{k,3}$ independent standard uniform variables
     $l_{k}\leftarrow F^{-1}_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}(w_{k,1}\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$
     $a_{k}\leftarrow F^{-1}_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}(w_{k,2}\,|\,z,\bar{l}_{k},\bar{a}_{k-1})$
     if k = 0 then
         $v_{k,j}=w_{0,1}$
     else
         for $j\leftarrow 1,\ldots,k$ do $\triangleright$ $L_{j}$ dist. in survivors per (14)
              $v_{k,j}\leftarrow\frac{v_{k-1,j}-C(v_{k-1,j},q_{k-1,j};\Theta_{k,k-j})}{1-q_{k-1,j}}$
         end for
     end if
     $\xi_{k+1}\leftarrow w_{k,3}$
     for $j\leftarrow 0,\ldots,k$ do $\triangleright$ ‘Unweaving’ as per (15)
         $\xi_{k-j}\leftarrow h^{-1}(\xi_{k+1-j},v_{k,k-j};\Theta_{K-j})$
     end for
     $\widetilde{y}_{k+1}\leftarrow F_{\widetilde{Y}_{k+1}\mid\bar{A}_{k},Y_{k-1}=0}^{-1}(\xi_{0}\,|\,do(\bar{a}_{k}))$
     $y_{k+1}=I(\widetilde{y}_{k+1}>1)$
     if $y_{k+1}=1$ then
         Stop
     else $\triangleright$ Prepare $q$s for the next time step per (16)
         $q_{k,0}\leftarrow F_{\widetilde{Y}_{k+1}\mid\bar{A}_{k},Y_{k-1}=0}(\widetilde{Y}_{k+1}=1\,|\,do(\bar{a}_{k}))$
         for $j\leftarrow 1,\ldots k$ do
              $q_{k,j}\leftarrow h\big{\{}q_{k,j-1},v_{k,j-1};\Theta_{k,j-1}\big{\}}$
         end for
         $k\leftarrow k+1$
     end if
end while
return $(\bar{l}_{k},\bar{a}_{k},\widetilde{y}_{k+1},\bar{y}_{k+1})$
Crucially, our method avoids computing the full conditional distribution of $\widetilde{Y}_{k+1}\mid\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0$, but rather requiring just a quantile corresponding to our simulated value; this is central to the efficiency of our approach. (14) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$. In (15), we recursively apply $k+1$ inverse $h$-functions to obtain the marginal quantile of $F(\widetilde{Y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0)$. Finally, we simulate $\widetilde{Y}_{k+1}$ from the target marginal structural model via inversion. Before proceeding to the next time step $k+1$, we prepare the $q_{k,j}$ through $k+1$ recursive computations in (16). In total, this time step requires $53k+$ 5 functions, including the simulation of $L_{k}$ and $A_{k}$.
@@
The proof is given in Appendix C.
In the derivation and proof presented above, we define the latent variable $\widetilde{Y}$ as the incremental time to failure which naturally aligns with the Cox marginal structural model that we use as a working example. Importantly, a key strength of our approach lies in its generality: our proposed algorithm is fully agnostic to the choice of survival model and holds for any specification of the latent variable and its distribution. This is a fundamental result that enables the direct simulation of survival outcomes from any causal marginal structural model, whether continuous or discrete.
@@
To illustrate this point, consider the case of simulating a binary survival outcome from a discrete marginal structural model, such as that defined in (10). Similar to the framework in Fan et al. (2017), we assume that there exists a continuous latent random variable, $S_{k+1}$, whose distribution under the intervention $do(\bar{A}_{k})$ is $F_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}$, such that the binary survival indicator is obtained by dichotomizing $S_{k+1}$ at a threshold $C$. This formulation allows dependencies between $Y_{k+1}$ and past covariates to be modelled via the latent variable $S_{k+1}$, and the simulation procedure follows the algorithm described above, replacing $\widetilde{Y}_{k+1}$ by $S_{k+1}$. Given the marginal quantile $u_{k+1}$, inversion yields $s_{k+1}=F^{-1}_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)$ Finally, the binary outcome is obtained as $y_{k+1}=I(s_{k+1}>C)$.
@@
In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(3k+5)=3K^{2}/2+7K/2$ calculations.This implies that the computational complexity scales quadratically with $K$.